# Foundation Models for Brain Signals: A Critical Review of Current Progress and Future Directions

### v1

@@ 3.1 Search Strategy
We conducted a comprehensive search across web platforms such as Google Scholar, arXiv, DBLP, IEEEXplore, bioRxiv, and medRxiv to identify relevant research in journals, conferences, workshops, and preprints. We limited our search query to EEG "Foundation Model". We then removed duplicate instances and manually reviewed the title, abstract, and introduction sections to confirm relevant EEG-FMs by identifying phrases similar to "We developed an EEG foundation model…" and "The proposed approach forms the basis for an EEG foundation model…". We note that our search starts from the year 2021 – the year the term Foundation Model was introduced [8] – and includes studies that were published or archived on or before September 30th, 2024.
@@ 3.2 EEG-FM Summaries
Table 1 summarizes all ten EEG-FMs identified in our search. Although all the EEG-FMs share many commonalities, each FM is unique in its own right. In order to highlight the building blocks of each FM and their unique strenghts, below we summarize each FM considering several factors, such as the amount of training data (in channel-hours), model size (in terms of the number of trainable parameters), the types of EEG data (scalp EEG and/or iEEG), the way inputs are configured (raw time series, power spectra, or time-frequency representation), architectural components (convolutional and/or transformer blocks), the SSL tasks used for pretraining (masked reconstruction, auto-regressive modeling, and/or contrastive learning), and the evaluations performed.
@@
Neuro-GPT [link] 540k 79.53M Fixed multi-channel time series data Scalp EEG Encoder with both convolution and transformer layers and GPT-2 as the decoder Masked reconstruction (causally masked latent embeddings)
Brant [link] 277k 68M, 104M, 249M and 506M Variable multi-channel time series data intracranial EEG Two transformer encoders for time and space and a linear decoder Masked reconstruction
BIOT [link] 312k 3.3M Variable multi-channel spectral data Scalp EEG Linear transformer, encoder only architecture Contrastive learning
EEGFormer 540k N/A Multi-channel spectral data Scalp EEG A transformer encoder and a shallow transformer decoder Codebook-based reconstruction
LaBraM [link] 80k 5.8M, 46M and 369M Fixed multi-channel time series data Scalp EEG Convolutional temporal encoder and transformer encoder layers and a linear decoder. A seperate decoder for tokenization. Masked reconstruction (token-level)
@@
NeuroLM [link] 800k 250M, 500M and 1.7B Variable multi-channel time series data Scalp EEG Vector quantization for tokenization with convolutional temporal encoder and transformer spatial encoder Autoregressive reconstruction (token-level)
@@
BrainWave 877k N/A Variable multi-channel spectrogram data Scalp EEG and intracranial EEG Transformer encoder with channel attention and a lightweight decoder Masked reconstruction (whole spectrogram)
@@
BrainBERT [15]: As the first released EEG-FM, BrainBERT is relatively smaller than others, with 43.18M parameters and was trained using a modest dataset of 4.5k channel-hours of iEEG data. The inputs were represented as channel-wise spectrograms and a BERT[65]-type model was trained to predict masked patches for different types of spectrograms (such as STFT and Superlets). The model comprised a transformer encoder with a shallow decoder with two linear layers. Evaluations showed generalizability to unseen subjects and unseen electrode locations; however, the test data were from the same distribution as the training data. The evaluations also showed that performing linear probing on BrainBERT embeddings was as good as training supervised deep neural networks (DNN) from scratch for most of the evaluation tasks, which are focused on predicting brain-evoked responses for watching movies. Additionally, their evaluations showed that fine-tuning BrainBERT embeddings can reach DNN performance from scratch with as little as 15% of the training data for one of the tasks. A task-agnostic intrinsic dimensionality [66] (ID)-based analysis showed that the BrainBERT embeddings of different brain regions had different IDs, compared to a relatively constant distribution across electrodes in randomly initialized weights.
@@
Neuro-GPT [16]: This mid-size EEG-FM with 79.53M parameters was trained entirely using scalp EEG data from the full TUH corpus, including 540k channel-hours of clinical scalp EEG data. The model takes raw time series of the 22 EEG channels in the standard 10-20 layout as input and learns EEG representations using a combination of convolution and transformer layers. Those representations are then used as input to a GPT-2 [67] decoder which autoregressively predict the masked latents. It is noteworthy that the decoder has more parameters than the encoder in this setup compared which is not a common practice in other EEG-FMs. This model was evaluated only on EEG data from a BCI motor imagery task with four classes with a different channel configuration than the training data. However, the downstream data were transformed to the original input configuration using an inverse-forward approach[68]. The results showed that fine-tuning or linear probing the pretrained model performed better than models trained from scratch, including some EEG-specific fully-supervised DL approaches.
@@
Brant [17]: Brant is a relatively larger FM with 500M parameters and was trained using 277k channel-hours of iEEG data. However, its pretraining data were limited to a single dataset comprising 9 subjects. This model also takes raw EEG time series as input and is able to take inputs with different channel configurations. The model consisted of a temporal encoder that learns long-term temporal dependencies, a spatial encoder that learns spatial correlations, and a simple linear decoder. The spatial encoder used in this model to capture spatial relationships is a novel contribution compared to previous EEG-FMs. This model also has three scaled-down versions with of 68M, 104M, and 249M parameters, respectively, that are trained on the same data. These models were evaluated on short/long term signal forecasting, frequency phase forecasting, imputation, and seizure detection tasks.
@@
EEGFormer [19]: This model was also pretrained on the whole TUH corpus, which includes approximately 540k channel-hours of clinical scalp EEG data. It included a transformer encoder and a shallow transformer decoder and it was pretrained using a masked-reconstruction objective. The encoder generated latents of input EEG patches are used to train a vector quantizer to match neural codes generated by a neural codebook. The decoder then reconstructs the original EEG patches using these neural codes. Evaluations included several downstream tasks derived from the TUH corpus and an out-of-distribution (OOD) evaluation on a neonatal seizure detection task. Additional experiments also included interpretability analyses using using the learned codebook.
@@
NeuroLM [22]: This model was inspired by a previous EEG-FM, LaBraM. However, NeuroLM was trained on 10x more data and is one of the largest EEG-FMs with 1.7B parameters, along with 250M and 500M parameter versions. NeuroLM utilized a text-aligned neural tokernizer, which is trained using temporal and frequency reconstruction along with a text/EEG domain classifier that is trained adversarially. The neural tokenizer is similar to that of LaBraM, except for the text alignment component and also included temporal reconstruction in addition to frequency reconstruction. However, despite this novel contribution, the evaluations indicated that the model loses some performance on downstream tasks compared to LaBraM and other state-of-the-art models.
@@
BrainWave [24]: This model was pretrained using a large dataset of size 877k channel-hours, including both scalp EEG and iEEG. It includes a transformer encoder and a channel attention module that transform EEG spectrograms into latent representations, which are then decoded by a lightweight decoder. This encoder-decoder architecture was trained using a masked-reconstruction objective. BrainWave is one of the few EEG-FMs trained and evaluated on both scalp EEG and iEEG signals, demonstrating the benefits of joint pretraining over unimodal approaches. The model has been extensively evaluated under different settings, such as cross-subject, cross-hospital, cross-subtype and few-shot classification, showcasing the generalizability and robustness across various clinical tasks.
@@ 4.1 Preparation and Representation of Input Data
Most EEG-FMs (e.g., LaBraM, EEGFormer, Neuro-GPT, Mentality, BIOT, and NeuroLM) were trained only on scalp EEG data, although some (e.g., Brant and BrainBERT) were trained exclusively on iEEG data. Two EEG-FMs, FoME and BrainWave, were trained on both scalp and intracranial EEG data. A large portion of the training data in the models utilized scalp EEG was taken from the Temple University Hospital (TUH) EEG corpus [5], which has a total of 540k channel-hours of scalp EEG data. Other scalp EEG datasets used for pretraining include CHB-MIT[73] and various small datasets collected for research purposes. Some FMs utilized EEG data available within PSG datasets such as SHHS[74], Sleep-EDF [75], and CAP Sleep [76]. Additionally, public or private iEEG data accounted for large portions of training data in some models, such as Brant and BrainWave. The common publicly available iEEG datasets used for pretraining were MAYO [77], FNUSA [77], Brain TreeBank[78], and CCEP [79]. Training data volume ranged from 14k channel-hours in BrainBERT to 877k channel-hours in BrainWave. BrainWave utilized the most diverse set of pretraining data, including clinical scalp EEG datasets, sleep EEG datasets, various smaller-scale scalp EEG repositories, and several iEEG datasets. LaBraM and NeuroLM also utilized moderately diverse datasets for pretraining, although all of the datasets were scalp EEG datasets.
@@
Input representation: EEG is a spatio-temporal data modality, which can be denoted as $X\in\mathbb{R}^{C\times T}$ with $C$ channels and $T$ time points. In addition to the spatial and temporal information, spectral information is also useful in interpreting EEG data, which is used most often as expert-derived features in statistical ML models. Furthermore, time-frequency representations, such as spectrograms or wavelet transforms, are also used to preserve information on all three domains. All EEG-FMs used one or a combination of these representations as input.
The most widely used input representation is the multivariate time series data format, which was adopted by six out of the ten EEG-FMs. With the adoption of transformer architecture, the windows that are segmented from the original recording require further segmentation into smaller patches, also known as tokens. These patches or tokens can then be added with temporal and spatial information with positional embeddings, which allows the models to be trained using EEG data with different channel configurations. This approach was adopted by models such as Brant, NeuroLM, and FoME, which allowed training using datasets with a varying number of channels. Two models, BIOT and EEGFormer, utilized power-spectral data as input instead of the original time series patches, while two other models, Brant and FoME, added power-spectral data to complement the raw time series inputs. Some EEG-FMs, such as LaBraM, Neuro-GPT, and Mentality, standardized different EEG datasets with different channel configurations to a fixed set of input channels. Two other models, BrainBERT and BrainWave, utilized channel-wise time-frequency representations of 1-5 seconds of EEG data as input. Various experiments performed in BrainBERT evaluations utilized spectrograms generated using classical methods like Short-Term Fourier Transform [80] (STFT) as well as scalograms [81] generated using modern methods such as Superlets [82].
@@
TUEG [83] traintrain traintrain traintrain traintrain traintrain
TUAB (Abnormal/normal classification - binary) $^{\#} evaleval^{*}$ $^{*} evaleval^{*}$
TUAR (EEG artifact classification - multiclass) $eval^{*}$ traintrain
TUEP (Epilepsy classification - binary) traintrain
TUEV (EEG Events classification - multiclass) $^{\#} evaleval^{*}$ $^{*} evaleval^{*}$
TUSL (EEG slowing classification - multiclass) $eval^{*}$ traintrain $eval^{*}$
TUSZ (Seizure type classification - multiclass) $eval^{*}$ traintrain $train/eval^{*}$
CHB-MIT [73] (Peadiatric seizure detection - binary) $eval^{\#}$ traintrain $eval^{\#}$
Sleep-EDFx [75] (Sleep stage classification - multiclass) $eval^{*} traintrain$
Siena Scalp EEG [84] (Seizure classification - binary) traintrain traintrain traintrain
SHHS [85][74] (Sleep stage classification - multiclass) traintrain
PREST (Abnormal event detection - binary) traintrain
CAP Sleep [76] (Sleep stage classification - multiclass) traintrain
HMC [86] (Sleep stage classification - multiclass) $eval^{\#}$ traintrain
SRM [87] (Resting state EEG data) traintrain
Schizophrenia-81 (Schizophrenia classification - binary) traintrain
Stroke-50 [88] (Hand movement classification - binary) traintrain
PD-31 [89] (Parkinson disease classification - binary) traintrain
IowaDataset [90] (Unknown task/s) traintrain
UNMDataset [91] (Parkinson disease classification - binary) traintrain
AD-184 [92] (Alzheimer’s disease classification - binary) traintrain
@@
BCI Competition IV-1 [102] (Movement classification - binary) traintrain traintrain
Emobrain [103] (Emotion classification - multiclass) traintrain traintrain
Grasp/Lift EEG Challenge [104] (Grasp classification/regression) traintrain traintrain
Inria BCI Challenge [105] (Next letter classification - multiclass) traintrain traintrain
EEG Motor Imagery Dataset [106] (Classification - multiclass) traintrain traintrain traintrain
Visual Categorization EEG Data [107] (Classification - multiclass) traintrain traintrain
Resting State EEG Data [108] (Eye state classification - binary) traintrain traintrain
SEED Series [109, 110, 111] (Emotion classification - multiclass) $^{*} train/evaltrain/eval^{*}$ $train/eval^{*}$
SPIS Resting State Dataset [112] (CVS and HRT regression) traintrain traintrain
Brain-Invaders [113] (Target vs. non-target classification - binary) traintrain traintrain
Multiple data sources [114, 115, 116, 117, 118] (Unknown tasks) traintrain traintrain
@@
Private dataset collection without source (Unknown) traintrain
@@
CCEP [79] (Seizure onset zone localization - binary) traintrain

### v2

# EEG Foundation Models: A Critical Review of Current Progress and Future Directions

@@ 3.1 Search Strategy
We conducted a comprehensive search across various web platforms, including Google Scholar, arXiv, DBLP, IEEE Xplore, bioRxiv, and medRxiv, to identify relevant research in journals, conferences, workshops, and preprints. We limited our search query to EEG "Foundation Model". We then removed duplicate instances and manually reviewed the title, abstract, and introduction sections to confirm relevant EEG-FMs by identifying phrases similar to "We developed an EEG foundation model…" and "The proposed approach forms the basis for an EEG foundation model…". We note that our search starts from the year 2021 – the year the term Foundation Model was introduced [8] – and includes studies that were published or archived on or before September 30th, 2024.
@@ 3.2 EEG-FM Summaries
Table 1 summarizes all ten EEG-FMs identified in our search. Although all the EEG-FMs share many commonalities, each FM is unique in its own right. In order to highlight the building blocks of each FM and their unique strengths, below we summarize each FM considering several factors, such as the amount of training data (in channel-hours, calculated as the total recording duration multiplied by the number of EEG channels; see supplement for details), model size (in terms of the number of trainable parameters), the types of EEG data (scalp EEG and/or iEEG), the way inputs are configured (raw time series, power spectra, or time-frequency representation), architectural components (convolutional and/or transformer blocks), the SSL tasks used for pretraining (masked reconstruction, auto-regressive modeling, and/or contrastive learning), and the evaluations performed.
@@
Neuro-GPT [link] 541k 79.53M Fixed multi-channel time series data Scalp EEG Encoder with both convolution and transformer layers and GPT-2 as the decoder Masked reconstruction (causally masked latent embeddings)
Brant [link] 281k 68M, 104M, 249M and 506M Variable multi-channel time series data intracranial EEG Two transformer encoders for time and space and a linear decoder Masked reconstruction
BIOT [link] 312k 3.3M Variable multi-channel spectral data Scalp EEG Linear transformer, encoder-only architecture Contrastive learning
EEGFormer 541k N/A Multi-channel spectral data Scalp EEG A transformer encoder and a shallow transformer decoder Codebook-based reconstruction
LaBraM [link] 80k 5.8M, 46M and 369M Fixed multi-channel time series data Scalp EEG Convolutional temporal encoder and transformer encoder layers and a linear decoder. A separate decoder for tokenization. Masked reconstruction (token-level)
@@
NeuroLM [link] 546k 250M, 500M and 1.7B Variable multi-channel time series data Scalp EEG Vector quantization for tokenization with convolutional temporal encoder and transformer spatial encoder Autoregressive reconstruction (token-level)
@@
BrainWave 878k N/A Variable multi-channel spectrogram data Scalp EEG and intracranial EEG Transformer encoder with channel attention and a lightweight decoder Masked reconstruction (whole spectrogram)
@@
BrainBERT [15]: As the first released EEG-FM, BrainBERT is relatively smaller than others, with 43.18M parameters and was trained using a modest dataset of 4.5k channel-hours of iEEG data. The inputs were represented as channel-wise spectrograms and a BERT[65]-type model was trained to predict masked patches for different types of spectrograms such as Short-Time Fourier Transform [66](STFT) and Superlets [67]. The model comprised a transformer encoder with a shallow decoder with two linear layers. Evaluations showed generalizability to unseen subjects and unseen electrode locations; however, the test data were from the same distribution as the training data. The evaluations also showed that performing linear probing on BrainBERT embeddings was as good as training supervised deep neural networks (DNN) from scratch for most of the evaluation tasks, which are focused on predicting brain-evoked responses for watching movies. Additionally, their evaluations showed that fine-tuning BrainBERT embeddings can reach DNN performance from scratch with as little as 15% of the training data for one of the tasks. A task-agnostic intrinsic dimensionality [68] (ID)-based analysis showed that the BrainBERT embeddings of different brain regions had different IDs, compared to a relatively constant distribution across electrodes in randomly initialized weights.
@@
Neuro-GPT [16]: This mid-size EEG-FM with 79.53M parameters was trained entirely using scalp EEG data from the full TUH corpus, including 541k channel-hours of clinical scalp EEG data. The model takes raw time series of the 22 EEG channels in the standard 10-20 layout as input and learns EEG representations using a combination of convolution and transformer layers. Those representations are then used as input to a GPT-2 [69] decoder which autoregressively predict the masked latents. It is noteworthy that the decoder has more parameters than the encoder in this setup compared which is not a common practice in other EEG-FMs. This model was evaluated only on EEG data from a BCI motor imagery task with four classes with a different channel configuration than the training data. However, the downstream data were transformed to the original input configuration using an inverse-forward approach[70]. The results showed that fine-tuning or linear probing the pretrained model performed better than models trained from scratch, including some EEG-specific fully-supervised DL approaches.
@@
Brant [17]: Brant is a relatively larger FM with 500M parameters and was trained using 281k channel-hours of iEEG data. However, its pretraining data were limited to a single dataset comprising 9 subjects. This model also takes raw EEG time series as input and is able to take inputs with different channel configurations. The model consisted of a temporal encoder that learns long-term temporal dependencies, a spatial encoder that learns spatial correlations, and a simple linear decoder. The spatial encoder used in this model to capture spatial relationships is a novel contribution compared to previous EEG-FMs. This model also has three scaled-down versions with of 68M, 104M, and 249M parameters, respectively, that are trained on the same data. These models were evaluated on short/long term signal forecasting, frequency phase forecasting, imputation, and seizure detection tasks.
@@
EEGFormer [19]: This model was also pretrained on the whole TUH corpus, which includes approximately 541k channel-hours of clinical scalp EEG data. It included a transformer encoder and a shallow transformer decoder and it was pretrained using a masked-reconstruction objective. The encoder generated latents of input EEG patches are used to train a vector quantizer to match neural codes generated by a neural codebook. The decoder then reconstructs the original EEG patches using these neural codes. Evaluations included several downstream tasks derived from the TUH corpus and an out-of-distribution (OOD) evaluation on a neonatal seizure detection task. Additional experiments also included interpretability analyses using using the learned codebook.
@@
NeuroLM [22]: This model was inspired by a previous EEG-FM, LaBraM. However, NeuroLM was trained on 7x more data and is one of the largest EEG-FMs with 1.7B parameters, along with 250M and 500M parameter versions. NeuroLM utilized a text-aligned neural tokernizer, which is trained using temporal and frequency reconstruction along with a text/EEG domain classifier that is trained adversarially. The neural tokenizer is similar to that of LaBraM, except for the text alignment component and also included temporal reconstruction in addition to frequency reconstruction. However, despite this novel contribution, the evaluations indicated that the model loses some performance on downstream tasks compared to LaBraM and other state-of-the-art models.
@@
BrainWave [24]: This model was pretrained using a large dataset of size 878k channel-hours, including both scalp EEG and iEEG. It includes a transformer encoder and a channel attention module that transform EEG spectrograms into latent representations, which are then decoded by a lightweight decoder. This encoder-decoder architecture was trained using a masked-reconstruction objective. BrainWave is one of the few EEG-FMs trained and evaluated on both scalp EEG and iEEG signals, demonstrating the benefits of joint pretraining over unimodal approaches. The model has been extensively evaluated under different settings, such as cross-subject, cross-hospital, cross-subtype and few-shot classification, showcasing the generalizability and robustness across various clinical tasks.
@@ 4.1 Preparation and Representation of Input Data
Most EEG-FMs (e.g., LaBraM, EEGFormer, Neuro-GPT, Mentality, BIOT, and NeuroLM) were trained only on scalp EEG data, although some (e.g., Brant and BrainBERT) were trained exclusively on iEEG data. Two EEG-FMs, FoME and BrainWave, were trained on both scalp and intracranial EEG data. A large portion of the training data in the models utilized scalp EEG was taken from the Temple University Hospital (TUH) EEG corpus [5], which has a total of 541k channel-hours of scalp EEG data. Other scalp EEG datasets used for pretraining include CHB-MIT[75] and various small datasets collected for research purposes. Some FMs utilized EEG data available within PSG datasets such as SHHS[76], Sleep-EDF [77], and CAP Sleep [78]. Additionally, public or private iEEG data accounted for large portions of training data in some models, such as Brant and BrainWave. The common publicly available iEEG datasets used for pretraining were MAYO [79], FNUSA [79], Brain TreeBank[80], and CCEP [81]. Training data volume ranged from 14k channel-hours in BrainBERT to 878k channel-hours in BrainWave. BrainWave utilized the most diverse set of pretraining data, including clinical scalp EEG datasets, sleep EEG datasets, various smaller-scale scalp EEG repositories, and several iEEG datasets. LaBraM and NeuroLM also utilized moderately diverse datasets for pretraining, although all of the datasets were scalp EEG datasets.
@@
Input representation: EEG is a spatio-temporal data modality, which can be denoted as $X\in\mathbb{R}^{C\times T}$ with $C$ channels and $T$ time points. In addition to the spatial and temporal information, spectral information is also useful in interpreting EEG data, which is used most often as expert-derived features in statistical ML models. Furthermore, time-frequency representations, such as spectrograms or wavelet [82] transforms, are also used to preserve information on all three domains. All EEG-FMs used one or a combination of these representations as input.
The most widely used input representation is the multivariate time series data format, which was adopted by six out of the ten EEG-FMs. With the adoption of transformer architecture, the windows that are segmented from the original recording require further segmentation into smaller patches, also known as tokens. These patches or tokens can then be added with temporal and spatial information with positional embeddings, which allows the models to be trained using EEG data with different channel configurations. This approach was adopted by models such as Brant, NeuroLM, and FoME, which allowed training using datasets with a varying number of channels. Two models, BIOT and EEGFormer, utilized power-spectral data as input instead of the original time series patches, while two other models, Brant and FoME, added power-spectral data to complement the raw time series inputs. Some EEG-FMs, such as LaBraM, Neuro-GPT, and Mentality, standardized different EEG datasets with different channel configurations to a fixed set of input channels. Two other models, BrainBERT and BrainWave, utilized channel-wise time-frequency representations of 1-5 seconds of EEG data as input. Various experiments performed in BrainBERT evaluations utilized spectrograms generated using classical methods like STFT as well as scalograms [82] generated using modern methods such as Superlets.
@@
TUEG [83] $train$ $train$ $train$ $train$ $train$
TUAB (Abnormal/normal classification - binary) $eval^{\#}$ $eval^{*}$ $eval^{*}$ $eval^{*}$
TUAR (EEG artifact classification - multiclass) $eval^{*}$ $train$
TUEP (Epilepsy classification - binary) $train$
TUEV (EEG Events classification - multiclass) $eval^{\#}$ $eval^{*}$ $eval^{*}$ $eval^{*}$
TUSL (EEG slowing classification - multiclass) $eval^{*}$ $train$ $eval^{*}$
TUSZ (Seizure type classification - multiclass) $eval^{*}$ $train$ $train/eval^{*}$
CHB-MIT [75] (Peadiatric seizure detection - binary) $eval^{\#}$ $train$ $eval^{\#}$
Sleep-EDFx [77] (Sleep stage classification - multiclass) $train/eval^{*}$ $train$
Siena Scalp EEG [84] (Seizure classification - binary) $train$ $train$ $train$
SHHS [85][76] (Sleep stage classification - multiclass) $train$
PREST (Abnormal event detection - binary) $train$
CAP Sleep [78] (Sleep stage classification - multiclass) $train$
HMC [86] (Sleep stage classification - multiclass) $eval^{\#}$ $train$
SRM [87] (Resting state EEG data) $train$
Schizophrenia-81 (Schizophrenia classification - binary) $train$
Stroke-50 [88] (Hand movement classification - binary) $train$
PD-31 [89] (Parkinson disease classification - binary) $train$
IowaDataset [90] (Unknown task/s) $train$
UNMDataset [91] (Parkinson disease classification - binary) $train$
AD-184 [92] (Alzheimer’s disease classification - binary) $train$
@@
BCI Competition IV-1 [102] (Movement classification - binary) $train$ $train$
Emobrain [103] (Emotion classification - multiclass) $train$ $train$
Grasp/Lift EEG Challenge [104] (Grasp classification/regression) $train$ $train$
Inria BCI Challenge [105] (Next letter classification - multiclass) $train$ $train$
EEG Motor Imagery Dataset [106] (Classification - multiclass) $train$ $train$ $train$
Visual Categorization EEG Data [107] (Classification - multiclass) $train$ $train$
Resting State EEG Data  [108] (Eye state classification - binary) $train$ $train$
SEED Series [109, 110, 111] (Emotion classification - multiclass) $train/eval^{*}$ $train/eval^{*}$ $train/eval^{*}$
SPIS Resting State Dataset [112] (CVS and HRT regression) $train$ $train$
Brain-Invaders [113] (Target vs. non-target classification - binary) $train$ $train$
Multiple data sources [114, 115, 116, 117, 118] (Unknown tasks) $train$ $train$
@@
Private dataset collection without source (Unknown) $train$
@@
CCEP [81] (Seizure onset zone localization - binary) $train$